OpenStudy (anonymous):

2 I-xI+3(-x) Please explain this to me and solve

OpenStudy (anonymous):

solve x or l

OpenStudy (anonymous):

The I was supposed to be absolute value brackets.. Sorry.. I think we're solving for x...

OpenStudy (anonymous):

ok is it \[2*\left| -x \right| +3*(-x) =0 ?\]

OpenStudy (anonymous):

Yes, just like that

OpenStudy (anonymous):

Given f(x)=2IxI+3x Find the following f(2-h)

OpenStudy (anonymous):

okay so we need to get the absolute value on one side, isolated so lets do that first:\[2\times|-x|+3\times(-x)=0\]\[2\times|-x|-3x=0\]\[2\times|-x|=3x\]\[|-x|={3x\over2}\]so now the answer is \[-x={3x\over2}~~~\implies~~~x=-{3x\over2}\]or\[{ -x=-{3x\over2}~~~\implies~~~x={3x\over2}}\]

OpenStudy (anonymous):

if \(f(x)=2|x|+3x\) the we just replace \(x\) with \(2-h\) to find \(f(2-h)\)...like so:\[f(2-h)=2|2-h|+3(2-h)\]\[f(2-h)=2|2-h|+6-3h~~~\implies~~~f(2-h)=2|2-h|-3h+6\]

OpenStudy (anonymous):

hope that helps @Precalc_sucks \(\large\ddot\smile\)

OpenStudy (anonymous):

You are the absolute best! Thank you so much for helping! Seeing the way you did the problem helped me make sense of it.

OpenStudy (anonymous):

YAY :D